OpenStudy (lolipop_drum312):

A line has slope −34 and y–intercept 5. Which answer is the equation of the line? Loading............

OpenStudy (lolipop_drum312):

A. y=−5x+3/4

OpenStudy (lolipop_drum312):

B. y=5x−3/4

OpenStudy (lolipop_drum312):

D. y=3/4x−5

OpenStudy (lolipop_drum312):

C. y=−3/4x+5

OpenStudy (lolipop_drum312):

Hmm? whatchu doin

OpenStudy (nerdychick16):

@wwhitlock

OpenStudy (nerdychick16):

but I think its B

OpenStudy (lolipop_drum312):

thats what i think but not sure

OpenStudy (nerdychick16):

@quickstudent

OpenStudy (nerdychick16):

@GIitter

OpenStudy (nerdychick16):

@awesomedude2078

OpenStudy (lolipop_drum312):

do u know glitter

OpenStudy (lolipop_drum312):

Glitter u there?

OpenStudy (lolipop_drum312):

oh yeah and that tops suppose to be -3/4

OpenStudy (giitter):

Im here

OpenStudy (giitter):

I not sure

OpenStudy (wwhitlock):

wolframalpha.com says y = 5-34 x